# Hybrid data assimilation techniques using the adjoint method in a coupled Lorenz system

### v1

@@ 3.1 SFDA (setup 1)
(a) Percentage error.
(b) Percentage uncertainty.
@@ 3.2 HDA (setup 2)
As the uncertainty of the parameters is independent of knowledge of the true parameters, this post-fit observable can be plotted for all climate models. Thus, we propose a quantitative procedure for verifying the optimal coupling strength, beyond simply observing synchronisation in the time series, in variational data assimilation. This procedure involves finding point of minimum parametric uncertainty in the linear region of the graph. It occurs at the first point the system is successfully synchronised, in this case $\alpha=7.5$. We acknowledge that such a procedure is only applicable on models with lower numbers of parameters where the global minimum of the cost function has been successfully found and running an assimilation multiple times is not adversely computationally costly. Running pre-fit scans of the cost function against a varying parameter would provide a more computationally efficient alternative to check if the cost function is smooth with no local minima for a fixed value of $\alpha$.
@@
(a) Percentage error.
(b) Percentage uncertainty.
@@
4 Conclusion
In this paper we have demonstrated the ability to constrain a Lorenz ’63 model using a second model with an adjoint by 4D-Var data assimilation. Such an approach removes the need to generate an adjoint for a forward model, if such an adjoint already exists for a separate, yet similar system. An important application of this technique in Earth system modelling would be a situation where a low-resolution ESM with an adjoint shares a parametrisation with a high-resolution ESM without an adjoint. The low-resolution ESM could provide all necessary variables, through synchronisation, to run its tangent linear adjoint model. This can then be utilised to estimate parameters in the high-resolution ESM without an adjoint. Moreover, using a simpler lower resolution version could make data assimilation quicker and would use fewer computational resources than an adjoint of the high-resolution ESM. We have also shown that running a forward model twice before beginning data assimilation can act to smooth the data and reduce the parametric uncertainty by roughly one third compared to a single forward run. Our overarching attention in this approach is optimising the parameters of a full Earth System model which will be tested as a next step. Nevertheless, it would also be possible to optimise the state variables which are more applicable to weather forecasting techniques. Future work will evaluate the resilience of hybrid data assimilation to mismodelling in the lower resolution model with an adjoint and examine how all setups in this research behave when assimilating sparse observations.
Acknowledgements

### v3

@@ 2.1 SFDA (setup 1)
((a)) Percentage error.
((b)) Percentage uncertainty.
@@ 2.2 HDA (setup 2)
((a)) Percentage error.
((b)) Percentage uncertainty.
@@
2.3 Mismodelling in HDA
HDA (setup 2) uses different forward and adjoint models with some shared physics. Therefore, it is important to examine the impact of a difference in the physics of these two models. We construct a test case where the equations of model 2, which has an adjoint, Eq. (6) are modified to give an oscillatory difference to both the truth and model 1. This can be done in a number of ways. We chose to introduce a multiplicative sine function to the equations in such a way it is also included in the adjoint matrix and thus modifies the gradient values returned to the fitting algorithm. Model 2 with an adjoint is now
$\frac{dx_{a}}{dt}$ $=$ $\sigma(y_{a}-x_{a})+\alpha(x_{f}-x_{a}),$ (16a)
$\frac{dy_{a}}{dt}$ $=$ $\rho x_{a}-y_{a}-x_{a}z_{a}+\alpha(y_{f}-y_{a}),$ (16b)
$\frac{dz_{a}}{dt}$ $=$ $x_{a}y_{a}-\beta z_{a}\cdot\left(1-\epsilon\sin\left(2\pi t\right)\right)$ (16c)
where $\epsilon$ is term which determines the strength of the oscillation term and $t$ is the simulation time. The effect of this term on the attractor is shown in Fig. 10, without synchronisation and assimilation. When compared to Fig. 2(a), it can be seen that this term is very effective at distorting the shape and probability density of the attractor.
@@ 2.3 Mismodelling in HDA
Figure 10: The Lorenz truth and model 2 attractors in the case of $\epsilon=1.0$. The bottom left and top right quadrants shows the attractors from all possible co-ordinate orientations. The diagonal plots show kernel density estimations (KDEs). No noise is added.
The consequence of varying this term on the accuracy of parameter optimisation after assimilation are shown in Fig. 11. With increasing epsilon the percentage error and uncertainty between fitted and truth systems remains stable. In spite of the large impact this term has on the attractor shape, this demonstrates a resilience of HDA to modelling differences between the forward-only model 1 and model 2 with an adjoint.
Figure 11: The percentage error (left) and uncertainty (right) between the true values of $(\sigma,\rho,\beta)$ and those from HDA. An ensemble of 100 assimilations is carried out over 100 different data sets. The median (line) and $68\%$ percentile intervals (shaded areas) are plotted. The noise level $25\%$ and, $\alpha=7.5$.
[…]
In this paper we have demonstrated the ability to constrain a Lorenz ’63 model using a second model with an adjoint by 4D-Var data assimilation. Such an approach removes the need to generate an adjoint for a forward model, if such an adjoint already exists for a separate, yet similar system. An important application of this technique in Earth system modelling would be a situation where a low-resolution ESM with an adjoint shares a parametrisation with a high-resolution ESM without an adjoint. The low-resolution ESM could, through synchronisation, provide all necessary variables to run its tangent linear adjoint model. This can then be utilised to estimate parameters in the high-resolution ESM without an adjoint. Moreover, using a simpler lower resolution version could make data assimilation quicker and would use fewer computational resources than an adjoint of the high-resolution ESM. We have also shown that running a forward model twice before beginning data assimilation can act to smooth the data and reduce the parametric uncertainty by roughly one third compared to a single forward run. Our overarching attention in this approach is optimising the parameters of a full ESM which will be tested as a next step. Nevertheless, it would also be possible to optimise the state variables which are more applicable to weather forecasting techniques. Future work will examine the resilience of such setups to spacially and temporally sparse data.
[…]
The pseudo-data samples used in this study are available on request from the authors.
[…]
PDK carried out the research, wrote the initial manuscript draft, and edited it. AB also assisted with the research and edited the manuscript. AK and DS designed the research concept and edited the manuscript.
[…]
The authors declare no potential competing interest.
Acknowledgements.
The figures in this research were plotted by the seaborn plotting package (Waskom, 2021) utilising our output which was managed using pandas (pandas development team, 2020). Automatic differentiation of our model was carried out using JAX (Bradbury et al., 2018). The minimisation of our cost function and uncertainty evaluation was done by iminuit (Dembinski and et al., 2020). This research was supported, in part, through the Koselleck grant $\textit{Earth}^{\textit{RA}}$ from the Deutsche Forschungsgemeinschaft (grant number: 66492934.) Contribution to the Centrum für Erdsystemforschung und Nachhaltigkeit (CEN) of Universität Hamburg. This work used resources of the Deutsches Klimarechenzentrum (DKRZ) granted by its Scientific Steering Committee (WLA).